# Thermodynamic topology of Black Holes in $F(R)$-Euler-Heisenberg gravity’s Rainbow

### v1

@@ IV Thermodynamics
Next, to ascertain the Hawking temperature, one has to consider primarily the surface gravity [106], which is supplied out of
@@
Exploring the total mass of black holes in $F(R)$-EH theory would be achievable by looking at the Ashtekar-Magnon-Das (AMD) approach [107, 108], which can be expressed as follows
@@ V Thermodynamic topology and photon sphere
Thermodynamic topology involves studying the thermodynamic properties of black holes using topological methods. It helps understand the critical points and phase transitions of black holes by analyzing their topological charges and numbers. A photon sphere is a region where gravity is strong enough that photons (light particles) are forced to travel in orbits. For a black hole, this sphere lies just outside the event horizon. The existence and properties of photon spheres can be studied using topological methods, which assign topological charges to these spheres. These charges help in understanding the stability and structure of the photon sphere. In recent studies, researchers have explored the relationship between thermodynamic topology and photon spheres in various black hole models, including those with hyperscaling violation. They have found that different topological charges correspond to different physical configurations, such as black holes and naked singularities[81]. Here, we want to study the Thermodynamic topology of Black Holes in $F(R)$-Euler-Heisenberg gravity’s Rainbow. so we follow some route that is explained in detail in the following section
@@ VI Thermodynamics topology of black holes
The primary concept associated with defects is the topological charge. To examine the thermodynamic topology of a black hole, we calculate the topological charge and utilize it to determine the topological classes. We employ Duan’s $\phi$ mapping technique to compute the topological charge. To introduce the thermodynamic properties of black holes, various quantities are utilized. For example, two distinct variables, such as mass and temperature, can describe the generalized free energy[109]. Given the relationship between mass and energy in black holes, we can reformulate our generalized free energy function as a standard thermodynamic function in the following form,
@@ VI.1 Topological classification of black holes in $F(R)$-Euler-Heisenberg gravity’s Rainbow
$\begin{split}&\mathcal{F}=-120(c+1)f_{\epsilon}g_{\epsilon}r_{H}^{7}+5(c+1)r_{H}^{6}\tau\left(r_{H}^{2}R_{0}-12g_{\epsilon}^{2}\right)\\ &+3f_{\epsilon}^{2}g_{\epsilon}^{2}q^{2}\tau\left(\lambda q^{2}-20r_{H}^{4}\right)\bigg{/}480\pi f_{\epsilon}g_{\epsilon}^{3}r_{H}^{5}\tau\end{split}$ (52)
@@
$\begin{split}&\phi^{r_{H}}=-16(c+1)f_{\epsilon}g_{\epsilon}r_{H}^{7}+(c+1)r_{H}^{6}\tau\left(4g_{\epsilon}^{2}-r_{H}^{2}R_{0}\right)\\ &+f_{\epsilon}^{2}g_{\epsilon}^{2}q^{2}\tau\left(\lambda q^{2}-4r_{H}^{4}\right)\bigg{/}32\pi f_{\epsilon}g_{\epsilon}^{3}r_{H}^{6}\tau\\ &\phi^{\Theta}=-\frac{\cot(\Theta)}{\sin(\Theta)}\\ \end{split}$ (53)
@@
$\begin{split}&\tau=\big{[}16(c+1)f_{\epsilon}g_{\epsilon}r_{H}^{7}\big{]}\bigg{/}\big{[}4cg_{\epsilon}^{2}r_{H}^{6}-cr_{H}^{8}R_{0}+f_{\epsilon}^{2}g_{\epsilon}^{2}\lambda q^{4}\\ &-4f_{\epsilon}^{2}g_{\epsilon}^{2}q^{2}r_{H}^{4}+4g_{\epsilon}^{2}r_{H}^{6}-r_{H}^{8}R_{0}{}\big{]}\end{split}$ (54)
We analyze Figure (3), which pertains to the structure of black holes in Euler-Heisenberg F(R)-Rainbow gravity. The figures display the normalized field lines. As depicted in the figures, some zero points represent the model’s topological charges with various free parameters. They are proportional to the winding number and are situated inside the blue contour at the coordinates $(r_{H},\Theta)$. We considered free parameters $q=0.5$, $\lambda=0.5$, $c=0.01$, $R_{0}=+0.001$ to draw these contours. In Figure (3,left), we plotted the curve for $f_{\epsilon}=g_{\epsilon}=1.2$. As shown in Figure (3,left), three blue loops represent the topological numbers $\omega$=(-1, +1, -1) and the total topological number $W=-1$. Also, in Figures (3,middle) and (3,right), we plotted the curve for $f_{\epsilon}=g_{\epsilon}=1.2,1.4$ respectively. (3,middle) represents the topological numbers $-1,+1$ with the total topological number $W=0$ and Figure (3,right) represents the one topological number $-1$, so that the total topological number is $W=-1$. Without loss of generality, we analyzed the topological properties of these black holes in this paper by choosing $\tau$ for black holes in Euler-Heisenberg F(R)-Rainbow gravity. The content discusses the stability of the black holes by examining the winding numbers and specific heat capacity. We have plotted Figures (4,top;left), (4,top; middle and righr) and (4,bottom; left and right) for the mentioned model with respect to free parameters $q=0.5$, $\lambda=0.5$, $c=0.01$, $R_{0}=-0.001$ and $f_{\epsilon}=g_{\epsilon}=1.2,1.4,1.6$ respectively, following the same approach. We plotted the normalized field vectors $n$ for these black holes in Figures (4,top;left), (4,top; middle and righr) and (4,bottom; left and right). Due to the better display of topological charges, we tried to draw two figures for each case, that is, Figure (4,top;left), Figures (4,top; middle and righr) together, and Figures (4,bottom; left and right) together are checked in the investigation of the total topological charges. As it is clear in these figures, only for Figure (4,top;left) corresponding to the free parameters $q=0.5$, $\lambda=0.5$, $c=0.01$, $R_{0}=-0.001$ and $f_{\epsilon}=g_{\epsilon}=1.2$, we have topological charges $\omega$=(-1, +1, -1) and the total topological charge $W=-1$. For the other two samples with $f_{\epsilon}=g_{\epsilon}=1.4,1.6$, the total topological charges are equal to $W=0$. By comparing Figures (3) and (4), it is clear that the positive or negative of $R_{0}$ and parameter values of $f_{\epsilon}=g_{\epsilon}$ play a very important role in determining the topological charges. We summarize the results in Table I.
@@
$q=0.5,\lambda=0.5,c=0.01,R_{0}=-0.001,f_{\epsilon}=g_{\epsilon}=1.4$     -1,+1,-1,+1             0
$q=0.5,\lambda=0.5,c=0.01,R_{0}=-0.001,f_{\epsilon}=g_{\epsilon}=1.6$     -1,+1,-1,+1             0
@@ VI.2 Photon sphere
In our past papers [96, 97], we demonstrated that the topological photon sphere can play a fundamental and reciprocal role in interpreting the behavior of a black hole. More precisely, since the existence of a photon sphere or photon rings is one of the essential requirements of ultra-compact gravitational structures, particularly black holes[98], the topological photon sphere can serve as a crucial test for determining the parameter ranges of black holes and, consequently, their overall behavior, and vice versa. To achieve this, we first construct the regular H potential and the vector field $\phi$ [99].
$\begin{split}H(r,\theta)=\sqrt{\frac{-g_{tt}}{g_{\varphi\varphi}}}=\frac{1}{\sin\theta}\bigg{(}\frac{f(r)}{h(r)}\bigg{)}^{1/2},\end{split}$ (55)
$\begin{split}&\phi^{r}=\frac{\partial_{r}H}{\sqrt{g_{rr}}}=\sqrt{g(r)}\partial_{r}H,\\ &\phi^{\theta}=\frac{\partial_{\theta}H}{\sqrt{g_{\theta\theta}}}=\frac{\partial_{\theta}H}{\sqrt{h(r)}},\end{split}$ (56)
which according to the length element and the following metric function:
[…]
$F(r)=1-\frac{m}{r}+\frac{f_{\varepsilon}^{2}q^{2}}{\left(1+c\right)r^{2}}+\frac{R_{0}r^{2}}{12g_{\varepsilon}^{2}}-\frac{\lambda q^{4}f_{\varepsilon}^{2}}{20r^{8}\left(1+c\right)},$ (58)
these functions will be as follows,
$H=\frac{\sqrt{900-\frac{900m}{r}+\frac{900f_{\varepsilon}^{2}q^{2}}{\left(1+c\right)r^{2}}+\frac{75R_{0}r^{2}}{g_{\varepsilon}^{2}}-\frac{45\lambda q^{4}f_{\varepsilon}^{2}}{r^{8}\left(1+c\right)}}}{30r\sin\!\left(\theta\right)},$ (59)
$\phi^{r}=\frac{3\csc\!\left(\theta\right)\left(\frac{2\left(-1-c\right)r^{8}}{3}+m\left(1+c\right)r^{7}-\frac{4f_{\varepsilon}^{2}q^{2}r^{6}}{3}+\frac{\lambda q^{4}f_{\varepsilon}^{2}}{6}\right)}{2r^{10}\left(1+c\right)},$ (60)
$\phi^{\theta}=-\frac{\sqrt{900-\frac{900m}{r}+\frac{900f_{\varepsilon}^{2}q^{2}}{\left(1+c\right)r^{2}}+\frac{75R_{0}r^{2}}{g_{\varepsilon}^{2}}-\frac{45\lambda q^{4}f_{\varepsilon}^{2}}{r^{8}\left(1+c\right)}}\,\cos\!\left(\theta\right)}{30r^{2}\sin\!\left(\theta\right)^{2}}.$ (61)
In this work, our objective is to investigate the impact of gravitational modifications on the effective parameters of the Euler-Heisenberg model within the context of Rainbow Gravity. Therefore, we specifically focus on the influence of parameters related to these structures. Also, given the limitation on the number of equations, we are compelled to consider some parameters as predetermined. An interesting point to note at the outset is that, as can be clearly seen from equation (60), the Rainbow parameter is effectively absent in this equation. Consequently, we will concentrate our efforts on the impact of the gravitational modification parameter.
VI.2.1 Case I: $f_{\varepsilon}=0.875<1$
@@
Figure 5: Metric function with different $\lambda$ for the black hole in F(R)-Euler-Heisenberg gravity’s Rainbow
According to the assumed values $m=1,q=0.5,c=0.01,R_{0}=0.001,g_{\varepsilon}=1.1$, we have drawn the metric function for different $\lambda$, in fig (5), the metric function has a root for $\lambda>-1.56$ and the black hole has an event horizon.
@@
Figure 6: The normal vector field $n$ in the $(r-\theta)$ plane. The photon spheres are located at $(r,\theta)=(1.173002072,1.57)$ with respect to $(\lambda=-1,q=0.5,m=1,c=0.01,R_{0}=0.001,g_{\varepsilon}=1.1)$, Right fig is the topological potential H(r) for black hole in F(R)-Euler-Heisenberg gravity’s Rainbow
Figure 7: The normal vector field $n$ in the $(r-\theta)$ plane. The photon spheres are located at $(r,\theta)=(0.9530605080,1.57)$, $(r,\theta)=(1.102139574,1.57)$ with respect to $(\lambda=-9,q=0.5,m=1,c=0.01,R_{0}=0.001,g_{\varepsilon}=1.1)$, Right fig is the topological potential H(r) for black hole in F(R)-Euler-Heisenberg gravity’s Rainbow
Euler-Heisenberg black holes Fix parametes Conditions *TTC  $(R_{PLPS})$
*Unauthorized area $q=0.5,m=1,c=0.01,R_{0}=0.001,g_{\varepsilon}=1.1$ $\lambda<-10.6168$ $nothing$  --
naked singularity $q=0.5,m=1,c=0.01,R_{0}=0.001,g_{\varepsilon}=1.1$ $-10.6168<\lambda<-1.54$ 00 --
unstable photon sphere $q=0.5,m=1,c=0.01,R_{0}=0.001,g_{\varepsilon}=1.1$ $\lambda\geq-1.54$ $-1$  $1.169958796$
Table 2: $\ast$ Unauthorized region: is the region where the roots of $\phi$ equations become negative or imaginary in this region. TTC: $\ast$ Total Topological Charge. $R_{PLPS}$: $\ast$ The minimum or maximum possible radius for the appearance of an unstable photon sphere.
@@
According to the method of determining topological charges [96, 97, 100] and can be inferred from Table 2, the chosen $\lambda$ in Figure (6) corresponds to a region where the system must possess an unstable photon sphere and a total topological charge of -1. This is confirmed by the potential function H(r) exhibiting a local maximum. Consequently, the structure will take the form of a normal black hole. However, in Figure (7), this selection pertains to a region classified as a naked singularity. As depicted in Figure (7), a global minimum appears in a space devoid of an event horizon, alongside a local maximum or the unstable photon sphere, resulting in a total topological charge of 0.
VI.2.2 Case II: $f_{\varepsilon}$=1
In the following, in order to avoid repetition, we will only mention the Figs and tables and finally we will discuss based on the tables. The metric function for different $\lambda$ ,The structure of topological charges and the diagram of the potential function H(r) of this state can be seen in Figure (8,9,10) and the parametric classification and general conclusions about different areas can be seen in Table 3.
@@ VI.2.2 Case II: $f_{\varepsilon}$=1
Figure 8: Metric function with different $\lambda$ for the black hole in F(R)-Euler-Heisenberg gravity’s Rainbow
@@
Figure 9: The normal vector field $n$ in the $(r-\theta)$ plane. The photon spheres are located at $(r,\theta)=(1.033146077,1.57)$ respect to with respect to $(\lambda=1,q=0.5,m=1,c=0.01,R_{0}=0.001,g_{\varepsilon}=1.1)$, Right fig is the topological potential H(r) for Euler-Heisenberg black hole model
Figure 10: The normal vector field $n$ in the $(r-\theta)$ plane. The photon spheres are located at $(r,\theta)=(0.7097591462,1.57)$, $(r,\theta)=(0.9921280053,1.57)$ with respect to $(\lambda=-0.544,q=0.5,m=1,c=0.01,R_{0}=0.001,g_{\varepsilon}=1.1)$, Right fig is the topological potential H(r) for black hole in F(R)-Euler-Heisenberg gravity’s Rainbow
Euler-Heisenberg black holes Fix parametes Conditions *TTC  $(R_{PLPS})$
*Unauthorized area $q=0.5,m=1,c=0.01,R_{0}=0.001,g_{\varepsilon}=1.1$ $\lambda<-1.544$ $nothing$  --
naked singularity $q=0.5,m=1,c=0.01,R_{0}=0.001,g_{\varepsilon}=1.1$ $-1.544<\lambda<-0.0135$ 00 --
unstable photon sphere $q=0.5,m=1,c=0.01,R_{0}=0.001,g_{\varepsilon}=1.1$ $\lambda\geq-0.0135$ $-1$  $1.009331770$
Table 3: *Unauthorized region: is the region where the roots of $\phi$ equations become negative or imaginary in this region. TTC: *Total Topological Charge. $R_{PLPS}$:* The minimum or maximum possible radius for the appearance of an unstable photon sphere .
VI.2.3 Case III: $f_{\varepsilon}$ = $1.5>1$
The structure of topological charges and the diagram of the potential function H(r) of this state can be seen in Figure (11) and the parametric classification and general conclusions about different areas can be seen in Table 4.
Figure 11: The normal vector field $n$ in the $(r-\theta)$ plane. The photon spheres are located at $(r,\theta)=(0.5644005040,1.57)$ with respect to with respect to $(\lambda=0.544,q=0.5,m=1,c=0.01,R_{0}=0.001,g_{\varepsilon}=1.1)$, Right fig is the topological potential H(r) for black hole in F(R)-Euler-Heisenberg gravity’s Rainbow
Euler-Heisenberg black holes Fix parametes Conditions *TTC
*Unauthorized area $q=0.5,m=1,c=0.01,R_{0}=0.001,g_{\varepsilon}=1.1$ $\lambda<0$ $nothing$
unstable photon sphere $q=0.5,m=1,c=0.01,R_{0}=0.001,g_{\varepsilon}=1.1$ $0<\lambda$ $-1$
Table 4: *Unauthorized region: is the region where the roots of $\phi$ equations become negative or imaginary in this region. TTC: *Total Topological Charge

### v2

@@ IV Thermodynamics
Next, to ascertain the Hawking temperature, one has to consider primarily the surface gravity Kubiznak:2016qmn , which is supplied out of
@@
Exploring the total mass of black holes in $F(R)$-EH theory would be achievable by looking at the Ashtekar-Magnon-Das (AMD) approach Ashtekar:1984zz ; Ashtekar:1999jx , which can be expressed as follows
@@ V Thermodynamic topology and photon sphere
Thermodynamic topology involves studying the thermodynamic properties of black holes using topological methods. It helps understand the critical points and phase transitions of black holes by analyzing their topological charges and numbers. A photon sphere is a region where gravity is strong enough that photons (light particles) are forced to travel in orbits. For a black hole, this sphere lies just outside the event horizon. The existence and properties of photon spheres can be studied using topological methods, which assign topological charges to these spheres. These charges help in understanding the stability and structure of the photon sphere. In recent studies, researchers have explored the relationship between thermodynamic topology and photon spheres in various black hole models, including those with hyperscaling violation. They have found that different topological charges correspond to different physical configurations, such as black holes and naked singularitiessa16 . Here, we want to study the Thermodynamic topology of Black Holes in $F(R)$-Euler-Heisenberg gravity’s Rainbow. so we follow some route that is explained in detail in the following section
@@ VI Thermodynamics topology of black holes
The primary concept associated with defects is the topological charge. To examine the thermodynamic topology of a black hole, we calculate the topological charge and utilize it to determine the topological classes. We employ Duan’s $\phi$ mapping technique to compute the topological charge. To introduce the thermodynamic properties of black holes, various quantities are utilized. For example, two distinct variables, such as mass and temperature, can describe the generalized free energyWei:2022dzw . Given the relationship between mass and energy in black holes, we can reformulate our generalized free energy function as a standard thermodynamic function in the following form,
@@ VI.1 Topological classification of black holes in $F(R)$-Euler-Heisenberg gravity’s Rainbow
$\begin{split}\mathcal{F}=\frac{-120(c+1)f_{\epsilon}g_{\epsilon}r_{H}^{7}+5(c+1)r_{H}^{6}\tau\left(r_{H}^{2}R_{0}-12g_{\epsilon}^{2}\right)+3f_{\epsilon}^{2}g_{\epsilon}^{2}q^{2}\tau\left(\lambda q^{2}-20r_{H}^{4}\right)}{480\pi f_{\epsilon}g_{\epsilon}^{3}r_{H}^{5}\tau}\end{split}$ (52)
@@
$\begin{split}&\phi^{r_{H}}=\frac{-16(c+1)f_{\epsilon}g_{\epsilon}r_{H}^{7}+(c+1)r_{H}^{6}\tau\left(4g_{\epsilon}^{2}-r_{H}^{2}R_{0}\right)+f_{\epsilon}^{2}g_{\epsilon}^{2}q^{2}\tau\left(\lambda q^{2}-4r_{H}^{4}\right)}{32\pi f_{\epsilon}g_{\epsilon}^{3}r_{H}^{6}\tau}\\ &\phi^{\Theta}=-\frac{\cot(\Theta)}{\sin(\Theta)}\\ \end{split}$ (53)
@@
$\begin{split}\tau=\frac{\big{[}16(c+1)f_{\epsilon}g_{\epsilon}r_{H}^{7}\big{]}}{\big{[}4cg_{\epsilon}^{2}r_{H}^{6}-cr_{H}^{8}R_{0}+f_{\epsilon}^{2}g_{\epsilon}^{2}\lambda q^{4}-4f_{\epsilon}^{2}g_{\epsilon}^{2}q^{2}r_{H}^{4}+4g_{\epsilon}^{2}r_{H}^{6}-r_{H}^{8}R_{0}{}\big{]}}\end{split}$ (54)
We analyze Figure (3), and (4), which pertain to the structure of black holes in Euler-Heisenberg F(R)-Rainbow gravity. The figures display the normalized field lines. As depicted in the figures, some zero points represent the model’s topological charges with various free parameters. They are proportional to the winding number and are situated inside the blue contour at the coordinates $(r_{H},\Theta)$. We have plotted Figures (3) for the mentioned model with respect to $q=0.5$, $\lambda=0.5$, $c=0.01$, $R_{0}=+0.001$ to draw these contours. In Figure (3), we plotted the curve for $f_{\epsilon}=g_{\epsilon}=1.2$. As shown in Figure (3), three blue loops represent the topological numbers $\omega$=(-1, +1, -1) and the total topological number $W=-1$. Also, in Figures (3) and (3), we plotted the curve for $f_{\epsilon}=g_{\epsilon}=1.2,1.4$ respectively. (3) represents the topological numbers $-1,+1$ with the total topological number $W=0$ and Figure (3) represents the one topological number $-1$, so that the total topological number is $W=-1$. Without loss of generality, we analyzed the topological properties of these black holes in this paper by choosing $\tau$ for black holes in Euler-Heisenberg F(R)-Rainbow gravity. The content discusses the stability of the black holes by examining the winding numbers and specific heat capacity. We have plotted Figures (4), for the mentioned model with respect to free parameters $q=0.5$, $\lambda=0.5$, $c=0.01$, $R_{0}=-0.001$ and $f_{\epsilon}=g_{\epsilon}=1.2,1.4,1.6$ respectively, following the same approach. We plotted the normalized field vectors $n$ for these black holes in Figures (4). As it is clear in these figures, for Figure (4) corresponding to the free parameters $q=0.5$, $\lambda=0.5$, $c=0.01$, $R_{0}=-0.001$ and $f_{\epsilon}=g_{\epsilon}=1.2$, we have topological charges $\omega$=(-1, +1, -1) and the total topological charge $W=-1$. For the other two samples with $f_{\epsilon}=g_{\epsilon}=1.4,1.6$, the total topological charges are equal to $\omega$=(-1, +1) and $\omega$=(-1) respectively which are shown in Figure (4) and Figure (4). By comparing Figures (3) and (4), it is clear that the positive or negative of $R_{0}$ and parameter values of $f_{\epsilon}=g_{\epsilon}$ play a very important role in determining the topological charges. We summarize the results in Table I.
@@
$q=0.5,\lambda=0.5,c=0.01,R_{0}=-0.001,f_{\epsilon}=g_{\epsilon}=1.4$        -1,+1             0
$q=0.5,\lambda=0.5,c=0.01,R_{0}=-0.001,f_{\epsilon}=g_{\epsilon}=1.6$           -1             -1
@@ VI.2 Photon sphere
In our past papers sa31 ; sa32 , we demonstrated that the topological photon sphere can play a fundamental and reciprocal role in interpreting the behavior of a black hole. More precisely, since the existence of a photon sphere or photon rings is one of the essential requirements of ultra-compact gravitational structures, particularly black holessa33 , the topological photon sphere can serve as a crucial test for determining the parameter ranges of black holes and, consequently, their overall behavior, and vice versa. To achieve this, we first construct the regular H potential and the vector field $\phi$ sa34 .
$\begin{split}H(r,\theta)=\sqrt{\frac{-g_{tt}}{g_{\varphi\varphi}}},\end{split}$ (55)
$\begin{split}&\phi^{r}=\frac{\partial_{r}H}{\sqrt{g_{rr}}},\\ &\phi^{\theta}=\frac{\partial_{\theta}H}{\sqrt{g_{\theta\theta}}},\end{split}$ (56)
So with respect to Eq. (22), we will have,
$H=\frac{\sqrt{36-\frac{36m}{r}-\frac{3r^{2}R_{0}}{g_{\varepsilon}^{2}}+\frac{36\left(\frac{q^{2}}{r^{2}}-\frac{q^{4}\lambda}{20r^{6}}\right)f_{\varepsilon}^{2}}{1+c}}\,g_{\varepsilon}}{6\sin\!\left(\theta\right)rf_{\varepsilon}},$ (57)
$\phi^{r}=\frac{3g_{\varepsilon}^{2}\csc\!\left(\theta\right)\left(\frac{2\left(-1-c\right)r^{6}}{3}+m\left(1+c\right)r^{5}-\frac{4q^{2}r^{4}f_{\varepsilon}^{2}}{3}+\frac{2q^{4}\lambda f_{\varepsilon}^{2}}{15}\right)}{2f_{\varepsilon}r^{8}\left(1+c\right)},$ (58)
$\phi^{\theta}=-\frac{\sqrt{36-\frac{36m}{r}-\frac{3r^{2}R_{0}}{g_{\varepsilon}^{2}}+\frac{36\left(\frac{q^{2}}{r^{2}}-\frac{q^{4}\lambda}{20r^{6}}\right)f_{\varepsilon}^{2}}{1+c}}\,g_{\varepsilon}^{2}\cos\!\left(\theta\right)}{6\sin\!\left(\theta\right)^{2}r^{2}f_{\varepsilon}}.$ (59)
Although our primary objective in this work is to examine the impact of gravitational corrections on the parameters of the Euler-Heisenberg model within the context of Rainbow Gravity, such as parameters $f_{\varepsilon}$, $g_{\varepsilon}$ , $c=f_{R_{0}}$ and $\lambda$, it appears that the parameter $R_{0}$ plays a more influential role. We will clearly observe that this parameter can affect the model’s behavior and essentially simulate the model and its associated photon sphere as either dS or AdS. Therefore, in this section, we will examine this parameter in both its negative and positive forms. In each case, we will separately analyze the impact of the applied corrections on the black hole action.
VI.2.1 Case I: $R_{0}<0$
In this scenario, according to the assumed values $m=1,R_{0}=-0.001,q=0.5,c=0.01,g_{\varepsilon}=1.1$ and $\lambda=0.5$ the model behaves entirely in the form of AdS, and only the event horizon appears, as shown in fig (5).
@@ VI.2.1 Case I: $R_{0}<0$
Figure 5: Metric function with different $f_{\varepsilon}$ for the black hole in F(R)-Euler-Heisenberg gravity’s Rainbow
Although we have examined different values for the existing parameters, but since they exhibit similar behavior, we will present one case for consistency with the values chosen in the previous sections.
@@
Figure 6: The normal vector field $n$ in the $(r-\theta)$ plane. The photon spheres are located at $(r,\theta)=(1.284148559,1.57)$ with respect to $(\lambda=0.5,q=0.5,m=1,c=0.01,R_{0}=-0.001,g_{\varepsilon}=1.1),f_{\varepsilon}=0.75$, Right fig is the topological potential H(r) for black hole in F(R)-Euler-Heisenberg gravity’s Rainbow
@@
According to the method of determining topological charges sa31 ; sa32 ; sa35 , the Figure (6) corresponds to a region where the system must possess an unstable photon sphere(U-PS) and a total topological charge of -1. This is confirmed by the potential function H(r) exhibiting a local maximum. Consequently, the structure will take the form of a normal black hole.Our investigations for various parameter values indicate that, in its AdS form and within the permissible parameter ranges, this black hole, unlike many previously studied black holes sa31 ; sa32 ; sa32.1 ; sa32.2 ; sa32.3 , does not exhibit the naked singularity region and consistently maintains the conditions of being a black hole for different parameter values.
VI.2.2 Case II: $R_{0}>0$
By selecting positive values for $R_{0}$, the model changes its behavior, exhibiting a dS form. In this scenario, typically two or three horizons can appear: the Cauchy horizon, the event horizon, and the cosmological horizon. Based on our chosen parameters $m=1,R_{0}=1,q=0.5,c=0.01,g_{\varepsilon}=1.1$ and $\lambda=0.5$, it appears that only the event horizon and cosmological horizon are observed, as shown in fig (7).
@@ VI.2.2 Case II: $R_{0}>0$
Figure 7: Metric function with different $f_{\varepsilon}$ for the black hole in F(R)-Euler-Heisenberg gravity’s Rainbow
Our previous studies on dS models sa31 ; sa32 ; sa32.1 ; sa32.2 ; sa32.3 indicate that due to the presence of the cosmological horizon, the model cannot exhibit any potential extremum beyond the cosmological horizon. In other words, it seems that no photon sphere model, whether stable or unstable, can emerge outside the cosmological horizon. Therefore, these photon spheres typically appear in the region between the event horizon and the cosmological horizon.
@@
Figure 8: The normal vector field $n$ in the $(r-\theta)$ plane. The photon spheres are located at $(r,\theta)=(1.282951427,1.57)$ respect to with respect to $(\lambda=0.5,q=0.5,m=1,c=0.01,R_{0}=1,g_{\varepsilon}=1.1,f_{\varepsilon}=0.75)$, Right fig is the topological potential H(r) for Euler-Heisenberg black hole model
In this context, the model, evaluated for multiple parameter values, consistently exhibits dS behavior and is unable to form a photon sphere outside the larger horizon or cosmological horizon for any parameter values. Consequently, as shown in fig (8), an unstable photon sphere with charge -1 always appears between the two horizons, corresponding to an energy maximum. It is important to note that in this case, we selected a value for $R_{0}$ that provided the best representation. For instance, with $R_{0}\simeq 0.001$, the cosmological horizon would appear at $r=190$, making its depiction in the figure somewhat challenging.